4. Determine the Ksp for X3Y2 if its solubility is 4.00 X 106 moles/Litre.
Answer #1

Calculation :

X3Y2 -------> 3X(2+) + 2Y(3-)

given solubility = 4 * 10^-6 mol /L

There is a 3:2 molar ratio between the X3Y2 that dissolves and X2+ that is in solution. In like manner, there is a 3:2 molar ratio between dissolved X3Y2 and Y3- in solution.

Ksp = [X2+]^3[Y3-]^2

Ksp = [4 * 10^-6]^3[4 * 10^-6]^2

Ksp = 1024 * 10^-30

Ksp = 1.02 * 10^-27

Answer

Ksp = 1.02 * 10^-27